# RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs

### v1

@@ 7 Conclusion
We present RiTeK, the first dataset specifically designed to evaluate the capability of models in handling complex reasoning over textual knowledge graphs (TKGs). This dataset offers diverse topological structures, relational types, entity types, and queries that integrate relational and textual information, requiring sophisticated reasoning across TKGs. RiTeK also includes rich textual descriptions for each node. To ensure the authenticity and accuracy of the queries, stringent validation was conducted by medical experts. In addition, to enhance the inference capabilities of LLMs, we are pioneering the adaptation of Monte Carlo Tree Search (MCTS) to textual knowledge graphs (TKGs) by introducing Relational MCTS, which retrieves relational path information along with relevant topological structures. Our extensive experiments on RiTeK reveal significant challenges for current models in handling both textual and complex relational information. Overall, RiTeK offers valuable opportunities for advancing research in complex reasoning tasks over textual knowledge graphs, while Relational MCTS provides a valuable way to retrieving complex relational paths from KGs in an unsupervised manner.
@@ 8 Limitations
The primary contribution of our work is the construction of two question-answering datasets based on textual knowledge graphs with complex topological structures. This work reveal significant challenges for current models in effectively handling both textual and complex relational information. We also demonstrated the potential of MCTS for complex reasoning path retrieval over textual knowledge graphs and validated the effectiveness of both MCTS and our proposed relational MCTS. However, the performance of relational MCTS in conjunction with other relational information retrieval methods on knowledge graphs remains unexplored, such as He et al. (2024); Baek et al. (2023) In the future, we will adapt these methods based on knowledge graphs to the context of textual knowledge graphs and investigate the effectiveness of these relational information retrieval methods.
@@ Appendix A Appendix
A.1 TKG resources
Ensembl  [^4] https://useast.ensembl.org/index.html,   [^5] https://www.nlm.nih.gov/research/umls/index.html, and Mondo Disease Ontology  [^6] https://mondo.monarchinitiative.org/.
A.2 Medical textual knowledge graph construction
@@
A.3 Data Analysis of query length and answer length
@@
Appendix B Relation MCTS
B.1 Textural Triplet Graph (TTG) and Input Question Encording
We begin our path retrieval approach by generating node embeddings in the TTG and creating a question embedding using a pre-trained language model. To elaborate, for each node $v$ in $\mathcal{G*}$, we utilize a pre-trained LLM, such as SentenceBert, to calculate the embedding of $v$, yielding the presentation $h_{v}$:
$h_{v}=LM(h:r:t:T(h):T(t))$
where, $T(*)$ is the textual description of head entity $h$ or tail entity $t$. The pre-trained LLM also is used to calculate the embedding of input question $q$, yielding the presentation $h_{q}$:
$h_{q}=LM(q)$
B.2 Relational Information Retrieving
Same as traditional MCTS, starting from the initial topic entity $e$, we perform multiple searches consisting of selection, expansion, simulations, and back-propagation. Specifically, In this approach, simulations are conducted using the default rollout policy, and to improve the accuracy of reward estimation, multiple rollouts are carried out. To maintain a balance between exploration and exploitation, we apply the Upper Confidence Bounds for Trees (UCT) method for node selection. The selection process is represented mathematically as:
$UCT(v,a)=\frac{Q(v,a)}{N(v,a)}+c\sqrt{\frac{lnN_{parent}(v)}{N(v,a)}}$
where $N(v,a)$ denotes the number of times node $v$ has been visited under action $a$ in prior iterations, and $N_{parent}(v)$ represents the visit count of the parent node of $v$. $Q(v,a)$ is the estimated reward value and will be updated through back-propagation. For each node, the reward function is computed based on the similarity between the query and a list that includes the current node and the subsequent node.
$Q(v,a)=Sim(h_{q},h_{v}^{1}:h_{v}^{2}:h_{v}^{d})$
where $d$ is the maximum depth of the searching tree. $c$ is a constant that balances exploitation and exploration. Once the search reaches a terminal node, either a terminal state (e.g. Stop Sign) or a predetermined maximum tree depth $d$, we obtain a trajectory from the root to terminal node.
Appendix C Relational Template
C.1 RiTeK-PharmKG

### v3

@@ 6 Conclusion
We present RiTeK, the first dataset specifically designed to evaluate the capability of models in handling complex reasoning over textual knowledge graphs (TKGs). This dataset offers diverse topological structures, relational types, entity types, and queries that integrate relational and textual information, requiring sophisticated reasoning across TKGs. RiTeK also includes rich textual descriptions for each node. To ensure the authenticity and accuracy of the queries, medical experts performed stringent validation. RiTeK sets a new standard for evaluating real-world retrieval systems. We evaluated 11 retrieval models on our benchmark dataset. Our experiments on RiTeK reveal significant challenges faced by current models in effectively handling both textual and relational information, especially under complex topological structures involving intricate relations and entities. RiTeK paves the way for future research aimed at advancing retrieval systems by emphasizing the need to enhance reasoning capabilities, particularly in retrieving complex reasoning paths under answer attribute constraints.
@@ 7 Limitations
RiTeK is currently limited to queries that involve only a single topic entity and rely solely on the textual and structural information present in the graph. Future work should explore the inclusion of multiple topic entities and incorporate additional modalities, such as images, to enable a more comprehensive and robust information retrieval system.
@@ Appendix A Appendix
A.1 Ethics Statement
All experiments in this study were conducted using publicly available datasets, including Prime, ADint, and PharmKG. These datasets contain only de-identified, non-personal, and non-sensitive information that was released for research purposes under appropriate data licenses. No private or confidential patient data were accessed or used. Consequently, no additional ethical approval was required. We employed AI tools to assist with grammar revision.
A.2 Hyperparameters
In the retrieval model, we set the maximum search depth to $d=3$ for both Random Walk and MCTS. The number of rollouts in MCTS is 16. For G-Retriever, KAR, GNN-RAG, and GCR, we adopted the same hyperparameters as reported in their respective source papers.
A.3 TKG resources
Ensembl  [^4] https://useast.ensembl.org/index.html, UMLS  [^5] https://www.nlm.nih.gov/research/umls/index.html, and Mondo Disease Ontology  [^6] https://mondo.monarchinitiative.org/.
A.4 Medical textual knowledge graph construction
@@
A.5 The prompt of Combining Textual and Relational Information
You are a creative assistant tasked with generating natural, diverse, and realistic queries by combining textual properties and relational templates. Write the query from the perspective of a <persona>, ensuring it is concise, human-like, and paraphrased while retaining the original meaning.
Consider the following characteristics for the persona:
  • Doctor: Formulate direct and practical questions aimed at diagnosing and treating. These questions should focus on side effects, symptoms, complications, and other clinically relevant aspects.
  • Medical Scientist: Generate detailed and specific questions reflecting the complexity of scientific inquiry. These questions should explore etiology, pathophysiology, genetic factors, pathways, proteins, or molecular functions.
  • Patient: Create straightforward questions that avoid professional medical terminology. These questions should focus on practical concerns, such as symptoms, effects, inheritance, or other relatable aspects, and may include more context from daily life.
Textual Properties: [<input_textual_properties>]
Relational Templates: [<input_relational_templates>]
Persona: <input_persona> (e.g., Doctor, Medical Scientist, Patient)
Ensure the query is realistic and diverse, leveraging flexibility in how the textual and relational elements are presented. Avoid directly copying the input phrases; instead, paraphrase them while retaining their original meaning. Please output only the generated query without any additional comments or explanations.
A.6 Data Analysis of query length and answer length
@@
Appendix B Relational Template
B.1 RiTeK-PharmKG
@@ B.2 RiTeK-ADint
Appendix C Topological Structures
C.1 Definition of Six Reasoning Topologies
Following the semantic structure framework introduced by Li and Ji (2022), we define six reasoning topologies that serve as the structural backbone for question generation and reasoning path simulation in RiTeK. Each topology represents a distinct reasoning pattern between entities in a textual knowledge graph (TKG), reflecting different levels of relational complexity and logical dependency.
  • 1-hop: The simplest reasoning structure, consisting of a single relation connecting the topic entity and the answer entity.
  • 2-hop: A two-step linear reasoning chain where the answer is connected to the topic entity through an intermediate entity.
  • 3-hop: A longer reasoning chain with three relational steps, representing more complex dependency and multi-level inference.
  • 1-hop with constraint: A single relational edge combined with a semantic or categorical constraint that filters valid answers.
  • 2-hop with constraint: A two-hop reasoning chain where the final answer is subject to an additional semantic or categorical restriction.
  • Two-to-one (Converging paths): Two distinct entities or relations converge on a common target entity. This topology reflects intersective reasoning, where the answer satisfies multiple relational constraints simultaneously.
C.2 Examples in the Medical Domain
  • 1-hop: For example, Amino Acid, Peptide, or Protein → [affects] → Disease or Syndrome illustrates a direct relationship such as “Which diseases are affected by a given protein?”
  • 2-hop: An example is Gene → [Interactions] → Chemical → [Treatment or therapy] → Disease, corresponding to the question “Which diseases can be treated by chemicals that interact with a specific gene?”
  • 3-hop: For instance, Gene → [Interactions] → Chemical → [Treatment or therapy] → Disease → [Biomarkers (diagnostic), or regulation linked to disease] → Gene models complex interactions such as “Which genes serve as biomarkers for diseases treated by chemicals interacting with a given gene?”
  • 1-hop with constraint: For example, Pharmacologic Substance → [treats] → Disease or Syndrome under the constraint “Disease type = neurodegenerative” represents questions like “Which pharmacologic substances treat neurodegenerative diseases?”
  • 2-hop with constraint: An example is Gene → [Enhance response, or activate, stimulate] → Gene → [Drug targets] → Disease with the constraint “Disease subtype = inflammatory”, corresponding to “Which genes activate other genes that target inflammatory diseases?”
  • Two-to-one (Converging paths): For instance, Gene 1 → [Interactions] → Chemical ← [Interactions] ← Gene 2 captures the question “Which chemicals interact with both Gene A and Gene B?”.
Appendix D Data Distribution Analysis
Shannon Entropy Type-Token Ratio
Medical domain
RiTeK-ADint 10.04 0.187
RiTeK-PharmKG 9.61 0.157
STARK-PRIME 9.63 0.143
General domain
STARK-AMAZON 10.39 0.179
STARK-MAG 10.25 0.180
Table 6: Query diversity measurement